Last year, Austin had
$20,000
to invest. He invested some of it in an account that paid
10%
simple interest per year, and he invested the rest in an account that paid
9%
simple interest per year. After one year, he received a total of
$1990
in interest. How much did he invest in each account?
First Account:
Second Account:

Answer by addingup(3677)

0.10x + 0.09(20,000-x) = 1990
0.10x + 1800 -0.09x = 1990
0.01x = 190
x = 19,000
OK, so 19,000 were invested at 10% and 1,000 at 9%
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
19,000(0.10) + 1,000(0.09) = 1990 Correct